Chemistry: Salt Solution. How much water must be evaporated from 32 ounces of a 4% salt solution to make a 6% salt solution?
Found 2 solutions by Alan3354, oberobic:
Answer by Alan3354(69443)

How much water must be evaporated from 32 ounces of a 4% salt solution to make a 6% salt solution?
-------------------
4% of 32 ounces is 1.28 oz of salt.
for 1.28 oz to be 6%, the total has to be 1.28/0.06 = 21.333...
32 - 21 1/3 = 10 2/3 ounces of water lost

Answer by oberobic(2304)

A key concept is view this type of problem in terms of the amount of 'pure stuff' you need.
.
You are told you have 32 oz of a 4% salt solution.
.
4%(32) = .04*32 = 1.28 oz of pure salt.
.
Given 1.28 oz of pure salt, how much 6% solution would that make?
.
Let 'x' be the unknown volume. Solve...
.
1.28/x = .06
.
Rearranging...
.
0.06x = 1.28
.
Multiply both sides by 100 to remove the decimals
.
6x = 128
.
Divide both sides by 6.
.
x = 21 1/3 = 21.3333
.
So a 21.3333 oz solution containing 1.28 oz of salt will be a 6% solution.
.
Looking back at the question, it asks how much must be evaporated.
.
32 - 21 1/3 = 10 2/3 = 10.667
.
Done.